OpenStudy (anonymous):

The vertices of a triangle are located at ,a(0, -8), B(- 4, 4 and C (2, 6) , and . If the triangle is dilated about the origin with a scale factor of 1/2 , what are the coordinates of ?c 1/3 2/3 1/6 4/12

OpenStudy (jdoe0001):

The vertices of a triangle are located at , , and . If the triangle is dilated about the origin with a scale factor of , what are the coordinates of ? you have one too many blanks, it might be better if you just take a quick screenshot of the material and post it here

OpenStudy (anonymous):

Basically, for each scale factor, just multiply each vertice's \(x\) and \(y\) values by the scale factor. For instance, let's call the scale factor \(k\) and the vertices are A(1,2), B(3,4), C(5,6). \[\textrm{Let }k = \frac{1}{3} \\ Ak = (1,2)*k = \left(\frac{1}{3}, \frac{2}{3}\right) \\ Bk = (3,4)*k = \left(1, \frac{4}{3}\right) \\ Ck = (5,6)*k = \left(\frac{5}{3}, 2\right) \]But we can't give you the full solutions if we don't know what your coordinates are.